OpenStudy (anonymous):

What amount of hydrogen ions, in moles, is present in 100 L of wine: [H+ (aq)] = 1 x 10^-3 mol/L, seawater: pH=8.00, and stomach acid: [H+(aq)]=10.0 mol/L

OpenStudy (anonymous):

amount of hydrogen ions in wine: \[10^{-3} mol/L * 100L = 0.1 mol\] amount of hydrogen ions in sea water : \[P_H = 8 \rightarrow -\log[H^+(aq)]= 8 \rightarrow [H^+(aq)]= 10^{-8} \] \[100L * 10^{-8} mlo/L = 10^{-6} mol \] and the amount of hydrogen ions in stomach acid is : \[100L*10mol/L = 1000mol\]